# Randomly punctured Reed–Solomon codes achieve list-decoding capacity over linear-sized fields

### v5

@@ 1 $B\leftarrow\emptyset$, $i_{0}\leftarrow\perp$, $\ell_{0}\leftarrow\perp$ for $\ell=1,\dots,j$ do
7             $B\leftarrow\emptyset$ for $\tau=1,\dots,2^{t}$ do
@@
            $M_{\ell}\leftarrow$ lexicographically smallest nonsingular $(t-1)k\times(t-1)k$ submatrix of $\mathsf{RIM}_{\mathcal{H},\mathcal{G}}^{B\cup\{i_{1},\dots,i_{\ell-1}\}}$ $\ell_{0}\leftarrow\ell$ // new refresh index
9
10
@@
2       $M_{1},\dots,M_{j}=\mathtt{GetMatrixSequenceRLC}(i_{1},\dots,i_{j-1})$ $i_{j}\leftarrow$ smallest index $i$ such that $M_{j}(X_{[i]\times[k]}=\alpha_{[i]\times[k]})$ is singular if $i_{j}$ not defined then
@@
$\mathop{\bf Pr\/}_{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\alpha_{[n]\times[k]}}}\left[{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\mathsf{RIM}_{\mathcal{H},\mathcal{G}}(X_{[n]\times[k]}=\alpha_{[n]\times[k]})}\text{ does not have full column rank}\right]\leq\binom{n}{r}2^{tr}\cdot\left({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\frac{t-1}{q}}\right)^{r}\ .$ (27)
We highlight that our probability bound here is better than the one in Lemma 3.1 for Reed–Solomon codes. This is because (i) all indeterminates in our generator matrix (and thus, the reduced intersection matrix) appear with degree 1 (rather than degree up to $k-1$), and (ii) our indeterminates are assigned independently uniformly at random, rather than random distinct values. Thus, the probability of any particular square submatrix matrix being made singular with an assignment is at most $\frac{t-1}{q}$, rather than $\frac{(t-1)k}{q-n}$: item (i) improves the numerator from $(t-1)k$ to $t-1$, and item (ii) improves the denominator from $q-n$ to $q$. This improved probability bound means we can use a smaller alphabet size for random linear codes than for Reed–Solomon codes. Other than this difference, the rest of our proof follows analogously. We include some more details for completeness.
@@
To prove Lemma 4.6, we similarly find a certificate $(i_{1},\dots,i_{r})$ for each singular reduced intersection matrix. This certificate is generated by an analogous algorithm, GetCertificateRLC, which uses an analogous helper function GetMatrixSequenceRLC. We show this certificate has the same three properties
@@
    There are few possible certificates
@@
Lemma 4.8 (Certificate count, Analogous to Corollary 3.10).
The number of possible outputs to $\mathtt{GetCertificateRLC}$ is at most $\binom{n}{r}2^{tr}$.
@@
Lemma 4.9 (Probability of one certficiate, Analogous to Corollary 3.12).
@@
$\mathop{\bf Pr\/}\left[{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\mathtt{GetCertificateRLC}(\alpha_{1,1},\dots,\alpha_{n,k})}=(i_{1},\dots,i_{r})\right]\leq\left({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\frac{t-1}{q}}\right)^{r}.$ (29)
@@
  • $E_{j}$ is the event that $M_{j}({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}X_{[i]\times[k]}=\alpha_{[i]\times[k]}})$ is non-singular for all $i<i_{j}$.
  • $F_{j}$ is the event that $M_{j}({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}X_{[i_{j}]\times[k]}=\alpha_{[i_{j}]\times[k]}})$ is singular.
The probability that all the events hold is at most $({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\frac{t-1}{q}})^{r}$.
@@
The proof is similar to the proof of Lemma 3.11. Lemma 3.11 follows from combining Eqaution (19) with the appropriate conditional probabilities. This lemma follows the same approach. We again assume without loss of generality $i_{1}<i_{2}<\cdots,i_{r}$.
@@
$\mathop{\bf Pr\/}_{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\alpha_{\{i_{j}\}\times[k]}}}\left[F_{j}|{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\alpha_{[i_{j}-1]\times[k]}}\right]\leq{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\frac{t-1}{q}}.$ (30)
To see (30), consider the determinant of $M_{j}({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}X_{[i_{j}-1]\times[k]}=\alpha_{[i_{j}-1]\times[k]}})$, a $(t-1)k\times(t-1)k$ matrix in $\mathbb{F}_{q}({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}X_{\{i_{j},i_{j}+1,\dots,n\}\times[k]}})$. View the determinant of $M_{j}({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}X_{[i_{j}-1]\times[k]}=\alpha_{[i_{j}-1]\times[k]}})$ as a polynomial in variables ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}X_{\{i_{j}+1,\dots,n\}\times[k]}}$ with coefficients in $\mathbb{F}_{q}[{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}X_{i_{j},1},\dots,X_{i_{j},k}}]$. It is nonzero because we assume $E_{j}$ holds, so there is some coefficient of the form $f({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}X_{i_{j},1},\dots,X_{i_{j},k}})$ that is nonzero. Since matrix $M_{j}$ has at most $t-1$ rows containing any variables among ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}X_{i_{j},1},\dots,X_{i_{j},k}}$, each appearing with total degree 1, the total degree of ${\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}X_{i_{j},1},\dots,X_{i_{j},k}}$ in the determinant of $M_{j}$ is at most $t-1$. Thus, the total degree of $f({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}X_{i_{j},1},\dots,X_{i_{j},k}})$ is at most $t-1$. Hence, by the Schwarz-Zippel lemma, $f$ becomes zero with probability at most $\frac{t-1}{q}$ over random $\alpha_{i_{j},1},\dots,\alpha_{i_{j},k}$. Thus, the probability that $F_{j}$ holds is at most $\frac{t-1}{q}$, giving (30).
@@
$\leq\mathop{\bf Pr\/}_{\alpha_{[n]\times[k]}}\left[\exists\text{ $(k+\varepsilon n)$-w.p.c. agreement hypergraph }\mathcal{H}\text{ such that }\mathsf{RIM}_{\mathcal{H},\mathcal{G}}({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}X_{[n]\times[k]}=\alpha_{[n]\times[k]}})\text{ not full column rank}\right]$
$\leq 2^{(L+2)n}\max_{\text{$(k+\varepsilon n)$-w.p.c. }\mathcal{H}}\quad\mathop{\bf Pr\/}_{\alpha_{[n]\times[k]}}\left[\mathsf{RIM}_{\mathcal{H},\mathcal{G}}({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}X_{[n]\times[k]}=\alpha_{[n]\times[k]}})\text{ not full column rank}\right]$
$\leq 2^{(L+2)n}\cdot\binom{n}{r}2^{(L+1)r}\left({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\frac{L}{q}}\right)^{r}\leq\left(2^{(L+2)n/r}\cdot\frac{en}{r}\cdot 2^{L+1}\cdot{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\frac{L}{q}}\right)^{r}\leq 2^{-Ln},$ (31)
as desired. Here, we used that $q={\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}2^{10L/\varepsilon}}$. ∎
4.4 Technical comparison with [GZ23]
To prove that random linear codes achieved list-decoding capacity (Theorem 1.3), we extended the framework for showing that (randomly punctured) Reed–Solomon codes achieve list-decoding capacity over linear-sized fields (Theorem 1.1). It is possible to instead use the framework of Guo and Zhang [GZ23] to show a similar result. However, using the framework of Guo and Zhang in the same way would have only worked for alphabet size that is linear in $n$, rather than, in our case, a near-optimal constant. Below, we explain why our new ideas were necessary for obtaining our near-optimal alphabet size.
In (31), our upper bound on the non-list-decodability probability is
@@ 4.4 Technical comparison with [GZ23]
where $r=\varepsilon n/2$, where $\varepsilon>0$ is roughly the gap to capacity. Here, the term $2^{(L+2)n}$ comes from a union bound over the number of possible hypergraphs, the term $\binom{n}{r}2^{(L+1)r}$ comes from a union bound over the number of possible certificates, and the term $\left(\frac{L}{q}\right)^{r}$ bounds the probability of a single certificate. We saw above that this probability is $o(1)$ as long as $q\geq 2^{10L/\varepsilon}$.
If we applied the framework of [GZ23] to random linear codes, the number of possible certificates would instead be $n^{r}$. Our bound on the non-list-decodability probability would then be
@@
For this to bound to be $o(1)$, we need to take $q\geq 2^{L/\varepsilon}\cdot n$, giving an alphabet size of $O(n)$. This would still have been a new result, as, previously, the Reed–Solomon codes of [GZ23] gave the smallest known alphabet size ($O(n^{2})$) of any linear code achieving list-decoding capacity with optimal list size $O(1/\varepsilon)$. However, using our framework allows us to achieve a near-optimal constant list size of $2^{O(L/\varepsilon)}$.
Acknowledgements
We thank Mary Wootters and Francisco Pernice for helpful discussions about [BGM23] and the hypergraph perspective of the list-decoding problem. We thank Karthik Chandrasekaran for helpful discussions about hypergraph connectivity notions and for the reference of Theorem 9 in [Fra11]. We thank Nikhil Shagrithaya and Jonathan Mosheiff for pointing out a mistake in the proof of Lemma 4.3 in an earlier version of the paper. We thank an anonymous reviewer for pointing out a mistake in the proof of Corollary A.4 in an earlier version of the paper.
Appendix A Alternate presentation of [BGM23]
Here, we include alternate presentations of some ideas from [BGM23]. Algebraically, our presentation is the same, but the hypergraph perspective streamlines combinatorial aspects of their ideas.

### v6

@@ 6    $i_{\ell-1}^{\prime}\leftarrow$ the $s$-th smallest element of $B$ that has type $\tau$
11    $B\leftarrow\emptyset$
12    for $\tau=1,\dots,2^{t}$ do
@@ 11    $B\leftarrow\emptyset$
14
15    $M_{\ell}\leftarrow$ lexicographically smallest nonsingular $(t-1)k\times(t-1)k$ submatrix of $\mathsf{RIM}_{\mathcal{H},\mathcal{G}}^{B\cup\{i_{1},\dots,i_{\ell-1}\}}$
    $\ell_{0}\leftarrow\ell$ // new refresh index
16
17
18
@@
2 $M_{1},\dots,M_{j}=\mathtt{GetMatrixSequenceRLC}(i_{1},\dots,i_{j-1})$
3 $i_{j}\leftarrow$ smallest index $i$ such that $M_{j}(X_{[i]\times[k]}=\alpha_{[i]\times[k]})$ is singular
4 if $i_{j}$ not defined then
@@ 2 $M_{1},\dots,M_{j}=\mathtt{GetMatrixSequenceRLC}(i_{1},\dots,i_{j-1})$
$\mathop{\bf Pr\/}_{\alpha_{[n]\times[k]}}\left[\mathsf{RIM}_{\mathcal{H},\mathcal{G}}(X_{[n]\times[k]}=\alpha_{[n]\times[k]})\text{ does not have full column rank}\right]\leq\binom{n}{r}2^{tr}\cdot\left(\frac{t-1}{q}\right)^{r}\ .$ (17)
We highlight that our probability bound here is better than the one in Lemma 3.1 for Reed–Solomon codes. This is because (i) all indeterminates in our generator matrix (and thus, the reduced intersection matrix) appear with degree 1 (rather than degree up to k1k-1), and (ii) our indeterminates are assigned independently uniformly at random, rather than random distinct values. Thus, the probability of any particular square submatrix matrix being made singular with an assignment is at most $\frac{t-1}{q}$, rather than $\frac{(t-1)k}{q-n}$: item (i) improves the numerator from (t1)k(t-1)k to t1t-1, and item (ii) improves the denominator from qnq-n to $q$. This improved probability bound means we can use a smaller alphabet size for random linear codes than for Reed–Solomon codes. Other than this difference, the rest of our proof follows analogously. We include some more details for completeness.
@@
To prove Lemma 4.6, we similarly find a certificate $(i_{1},\dots,i_{r})$ for each singular reduced intersection matrix. This certificate is generated by an analogous algorithm, GetCertificateRLC, which uses an analogous helper function GetMatrixSequenceRLC. We show this certificate has the same three properties:
@@
    There are few possible certificates.
@@
Lemma 4.8 (Analogous to Corollary 3.10).
The number of possible outputs of $\mathtt{GetCertificateRLC}$ is at most $\binom{n}{r}2^{tr}$.
@@
Lemma 4.9 (Probability of one certificate, analogous to Corollary 3.12).
@@
$\mathop{\bf Pr\/}\left[\mathtt{GetCertificateRLC}(\alpha_{1,1},\dots,\alpha_{n,k})=(i_{1},\dots,i_{r})\right]\leq\left(\frac{t-1}{q}\right)^{r}.$ (19)
@@
  • $E_{j}$ is the event that $M_{j}(X_{[i]\times[k]}=\alpha_{[i]\times[k]})$ is non-singular for all $i<i_{j}$.
  • $F_{j}$ is the event that $M_{j}(X_{[i_{j}]\times[k]}=\alpha_{[i_{j}]\times[k]})$ is singular.
The probability that all the events hold is at most $(\frac{t-1}{q})^{r}$.
@@
The proof is similar to the proof of Lemma 3.11. Lemma 3.11 follows from combining Equation (12) with the appropriate conditional probabilities. This lemma follows the same approach. We again assume without loss of generality $i_{1}<i_{2}<\cdots,i_{r}$.
@@
$\mathop{\bf Pr\/}_{\alpha_{\{i_{j}\}\times[k]}}\left[F_{j}|\alpha_{[i_{j}-1]\times[k]}\right]\leq\frac{t-1}{q}.$ (20)
To see (20), consider the determinant of $M_{j}(X_{[i_{j}-1]\times[k]}=\alpha_{[i_{j}-1]\times[k]})$, a $(t-1)k\times(t-1)k$ matrix in $\mathbb{F}_{q}(X_{\{i_{j},i_{j}+1,\dots,n\}\times[k]})$. View the determinant of $M_{j}(X_{[i_{j}-1]\times[k]}=\alpha_{[i_{j}-1]\times[k]})$ as a polynomial in variables $X_{\{i_{j}+1,\dots,n\}\times[k]}$ with coefficients in $\mathbb{F}_{q}[X_{i_{j},1},\dots,X_{i_{j},k}]$. It is nonzero because we assume $E_{j}$ holds, so there is some coefficient of the form $f(X_{i_{j},1},\dots,X_{i_{j},k})$ that is nonzero. Since matrix $M_{j}$ has at most t1t-1 rows containing any variables among $X_{i_{j},1},\dots,X_{i_{j},k}$, each appearing with total degree 1, the total degree of $X_{i_{j},1},\dots,X_{i_{j},k}$ in the determinant of $M_{j}$ is at most t1t-1. Thus, the total degree of $f(X_{i_{j},1},\dots,X_{i_{j},k})$ is at most t1t-1. Hence, by the Schwartz–Zippel lemma, $f$ becomes zero with probability at most $\frac{t-1}{q}$ over random $\alpha_{i_{j},1},\dots,\alpha_{i_{j},k}$. Thus, the probability that $F_{j}$ holds is at most $\frac{t-1}{q}$, giving (20).
@@
$\mathop{\bf Pr\/}_{\alpha_{[n]\times[k]}}\left[\text{Code generated by }\mathcal{G}|_{X_{[n]\times[k]}=\alpha_{[n]\times[k]}}\text{ is not $\left(\frac{L}{L+1}(1-R-\varepsilon),L\right)$ list-decodable}\right]$
@@
$\leq 2^{(L+2)n}\max_{\text{$(k+\varepsilon n)$-w.p.c. }\mathcal{H}}\quad\mathop{\bf Pr\/}_{\alpha_{[n]\times[k]}}\left[\mathsf{RIM}_{\mathcal{H},\mathcal{G}}(X_{[n]\times[k]}=\alpha_{[n]\times[k]})\text{ is not full column rank}\right]$
$\leq 2^{(L+2)n}\cdot\binom{n}{r}2^{(L+1)r}\left(\frac{L}{q}\right)^{r}\leq\left(2^{(L+2)n/r}\cdot\frac{en}{r}\cdot 2^{L+1}\cdot\frac{L}{q}\right)^{r}\leq 2^{-Ln},$ (21)
as desired. Here, we used that $q=2^{10L/\varepsilon}$. ∎
Remark 4.11.
Our tighter bound on the number of certificates that leveraged the symmetries of the matrices RIM is crucial to obtaining a near-optimal constant alphabet size for Theorem 1.3; we would only have obtained linear $O(n)$ alphabet size otherwise. In (21), our upper bound on the non-list-decodability probability is
@@
where $r=\varepsilon n/2$, where $\varepsilon>0$ is roughly the gap to capacity. Here $\binom{n}{r}2^{(L+1)r}$ is the number of possible certificates. If we had naively bounded the number of certificates by $n^{r}$, our bound on the non-list-decodability probability would then be
@@
For this bound to be $o(1)$, we need to take $q\geq 2^{L/\varepsilon}\cdot n$, giving an alphabet size of $O(n)$. This would still have been a new result, but leveraging the symmetries allowed us to achieve a near-optimal constant list size of $2^{O(L/\varepsilon)}$.
5 Alphabet Size Limitations
In this section, we establish Proposition 1.5. For positive integers $m$, view $\mathbb{F}_{2^{m}}$ as a vector space of dimension $m$ over base field $\mathbb{F}_{2}$. For a set $S\subseteq\mathbb{F}_{2^{m}}$, let
Lemma 5.1 (Proposition 3.2 of [3]).
Let $L$ be a subspace of $\mathbb{F}_{2^{m}}$. Then $P_{L}$ has the form
$X^{2^{\dim L}}+\sum_{i=0}^{\dim L-1}\alpha_{i}X^{2^{i}},$ (25)
where $\alpha_{i}\in\mathbb{F}_{2^{m}}$.
As an immediate corollary, we have
Lemma 5.2.
Let $L$ be an affine subspace of $\mathbb{F}^{2^{m}}$. Then $P_{L}$ has the form
$X^{2^{\dim L}}+\sum_{i=0}^{\dim L-1}\alpha_{i}X^{2^{i}}+\beta$ (26)
for $\alpha_{i},\beta\in\mathbb{F}_{2^{m}}$.
Proof.
Since $L$ is an affine subspace, there exists $\gamma$ such that $L-\gamma\stackrel{{\scriptstyle\rm def}}{{=}}\{\alpha-\gamma:\alpha\in L\}$ is a subspace of $\mathbb{F}_{2^{m}}$. By Lemma 5.1, we have $P_{L-\gamma}$ is of the form
$X^{2^{\dim L}}+\sum_{i=0}^{\dim L-1}\alpha_{i}X^{2^{i}}$ (27)
Lemma 5.3 (Analogous to Lemma 3.5 of [3]).
Proof.
so there are at least $2^{v(m-v)}$ affine-subspaces $L$ with $|L\cap S|\geq n/2^{m-v}$. For all such affine-subspaces $L$, the polynomial $P_{L}(X)$ has the form $X^{2^{v}}+\sum_{i=0}^{v-1}\alpha_{i}X^{2^{i}}+\beta$ by Lemma 5.2. Among these affine-subspaces $L$, by the pigeonhole principle, for at least a fraction $2^{-m(v-u-1)}$ of these subspaces, their subspace polynomials $P_{L}(X)$ have the same $\alpha_{i}$ for $i=u+1,u+2,\dots,v$. Let $\mathcal{L}$ be this family of subspaces. The number of subspaces is at least $2^{v(m-v)}\times 2^{-m(v-u-1)}=2^{(u+1)m-v^{2}}$, so $\mathcal{L}$ is the desired family of affine subspaces. ∎
Proof of Proposition 1.5.
Let $\delta=2^{-r-1}$ as in the statement of Proposition 1.5. Consider a Reed–Solomon code of length $n$ and rate $\delta$ over $\mathbb{F}_{q}$, where $q=2^{m}$ with $m$ sufficiently large. Let $S\subseteq\mathbb{F}_{q}$ be the set of $n$ evaluation points. Apple Lemma 5.3 with $u=m-\lceil{1.99r}\rceil$ and v=mrv=m-r. This gives a family $\mathcal{L}$ of $2^{m(m-\lceil{1.99r}\rceil)-(m-r)^{2}}=2^{2rm-\lceil{1.99r}\rceil m+r^{2}}\geq q^{\Omega(\log(1/\delta))}$ affine subspaces $L\leq\mathbb{F}_{2^{m}}$ for which $|L\cap S|\geq n/2^{m-v}=2\delta n$. Furthermore, for $L\in\mathcal{L}$, the subspace polynomials $P_{L}$ each have $2^{v}$ roots, and agree on all coefficients of degree larger than $2^{u}$. Let $L_{0}$ be an arbitrary element of $\mathcal{L}$. Then the polynomials $\{P_{L_{0}}-P_{L}:L\in\mathcal{L}\}$ are each of degree at most $2^{u}=2^{-\lceil{1.99r}\rceil}q\leq 4\delta^{1.99}q\leq\delta n$, and each agree with $P_{L_{0}}(X)$ on at least $|L\cap S|\geq 2\delta n$ values of $S$. Thus, our Reed–Solomon code is not $(1-2\delta,n^{\Omega(1/\delta)})$-list-decodable, as desired. ∎
Appendix
[…]
Here, we include alternate presentations of some ideas from [9]. Algebraically, our presentation is the same, but the hypergraph perspective streamlines combinatorial aspects of their ideas.